OpenStudy (anonymous):

**WILL MEDAL** What is the point-slope of the equation of a line that passes through thee point (6,8) and has a slope of -2? A. y+8=2(x-6) B. y+8=-2(x-6) C. y-6=-2(x-8) D. y-8=-2(x+6)

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

I AM HERE!

OpenStudy (cwrw238):

y-y1 = m(x-x1) where y1 = 8, x1=6 and m=-2

OpenStudy (cwrw238):

yes C is correct

OpenStudy (anonymous):

leave it to cwrw...

OpenStudy (anonymous):

Thx(:

OpenStudy (cwrw238):

oh hold on...

OpenStudy (cwrw238):

not correct is y-8 = -2(x-6)

OpenStudy (cwrw238):

none of them maybe b is written incorrectly

OpenStudy (cwrw238):

or C for that matter! lol

OpenStudy (anonymous):

Oh ok

OpenStudy (cwrw238):

C is incorrect because its got x1 = 8 and y1 = 6 and it should be the other way around